can you help me with the red marks
Answer #1

Data:    

n = 50   

μ = 15   

s = 6   

x-bar = 14.15   

Hypotheses:   

Ho: μ = 15   

Ha: μ ≠ 15   

Decision Rule:   

α = 0.05   

Lower Critical z- score = -1.959963985

Upper Critical z- score = 1.959963985

Reject Ho if |z| > 1.959963985

Test Statistic:   

SE = s/√n = 6/√50 = 0.848528137

z = (x-bar - μ)/SE = (14.15 - 15)/0.848528137423857 = -1.00

p- value = 0.3173

Decision (in terms of the hypotheses):

Since 1.001734607 < 1.959963985 we fail to reject Ho

Conclusion (in terms of the problem):

There is no sufficient evidence that μ is different from 15.

ANSWERS:  

(a) -1.00

(b) 0.3173

(c) No

(d) ± 1.96